THE ENERGY LEVELS OF URANIUM V

\underline{SCOTT R. BROZELL} and RUSSELL M. PITZER, Department of Chemical Physics, The Ohio State University, Columbus, OH 43210.

Spin-orbit configuration-interaction calculations are performed on U^{4+}. Uranium is modeled with relativistic effective core potentials and Gaussian cc-pVDZ and cc-pVTZ basis sets. This 5f^{2} ion has been studied extensively, and the calculations are compared with the experimental spectrum^{a}.

aJ. F. Wyart, V Kaufman, and J. Sugar Phys. Scr. 22, 389 (1980)